#20    We are interested in looking at the percentage of households that own dogs in...

#20

   We are interested in looking at the percentage of households that own dogs in the U.S. and England. We are given the following information.

Number of people surveyed    Number of people who own a dog

   U.S. 700    294

   England    850    391

The point estimate for the difference in proportions between people in the U.S who own dogs and people in England who own dogs is?
Answer #1

Given:

Country Number of people surveyed Number of people who own a dog
US 700 294
England 850 391

Proportion of people in the US who own dogs = 294/700 = 0.42

Proportion of people in the England who own dogs = 391/850 = 0.46

The difference in proportions between people in the US who own dogs and people in England who own dogs = 0.42 - 0.46 = -0.04